what is formula of projectile motion


1
Expert's answer
2021-08-02T08:41:52-0400

Given:

"v_0,\; \theta"

Equations of motion of a projectile are as follows

"x(t)=v_0\cos\theta* t\\
y(t)=v_0\sin\theta* t-\frac{gt^2}{2}"

The range

"R=\frac{v_0^2\sin2\theta}{g}"

The maximum height

"H=\frac{v_0^2\sin^2\theta}{2g}"